Project

General

Profile

NDOS Commissioning Meeting 2011/08/15

Attendance: Jon, Kurt, Minerba, Andrew, Mark Mathis, Jarek, Rafael, Leon

Agenda:

  1. Beam & Detector Update
  2. DAQ Status & Plans
  3. DCS Status & Plans
  4. AOB

Beam & NDOS Update

  • Lost beam over the weekend due to water leak in Linac. The leak was fixed sooner than expected and we got beam yesterday.
  • Strange behavior this morning. The run was still going from last night, but the RC GUI, Msg Viewer and RsrcMgr GUI had disappeared. Also, the background color of all nova-daq-[1-3] changed from bright blue to black. In the process of getting another run going, the RC GUI crashed again. Jon managed to track the problem down a bug in the NovaDatabase package, which has now been addressed.

DAQ Status & Plans

  • Leon:
    • wondering if any newly cooled APDs have gone bad. Jon: not yet, but it's been
    • Bill Gilbert made some progress on enabling the extended tests, CalTech should be ready to use this later today.
  • Denis:
    • took DSO scans for the latest 3 dcms where Denis and Leon installed ferrite beads on.
    • analysis is in progress, should complete the analysis today. So far results look good for both $0.10 and $0.40 beads. There is still some lingering noise on a few channels.
    • out of 50 FEBs, ~6 remain noisy (on the edge).
    • will present results at tomorrow's electronics meeting.
  • Andrew:
    • noticed on Friday that dcms were not being put into external timing mode thus some were not in sync (time) with each other, probably due to the fact that we're no longer loading firmware.
    • DAQ group will investigate the root cause, but in the meantime shifters are asked to force dcms into external timing mode prior to hitting the "Begin Run" button.
  • Denis & Andrew: we also need to understand why the beam spill server does not start reliably.

DCS Status & Plans

  • Nothing new to report.

AOB

  • None